zum Inhalt springen

News-Archiv

11.10.2019 - NMR-Abteilung: All spectrometers are back in service

Reminder: From next week the av300oc is switched into automation mode at 2pm.

On Monday and Tuesday 14th and 15th of…

9.10.2019 - NMR-Abteilung - Einschränkungen im Messbetrieb am 9.10. + 10.10. der OC - Update

Tomorrow, 10th of October, the measuring operation will be closed until construction work is finished and the floor is…

8.10.2019 - NMR-Abteilung: Limited access/possible closure of NMR labs on October 9 & 10

The NMR labs in organic chemistry may be under restricted access or closed due to planned construction work.

 

19.9.2019 - NMR-Abteilung: Maintenance work on av300oc on Thuesday 24th of September

Next Tuesday, av300oc will be temporarily out of service

16.9.2019 - NMR-Abteilung: Facility closed over weekend

Next weekend our facility will be closed, instruments are only accessibly until Friday at noon.

14.7.2019 - No access to NMR labs in organic chemistry

On Wednesday, 14th of august, there will be construction work in our corridor: Access to all NMRs in the organic…

18.7.2019 - av600 out of service

Dear users,

 

due to a total damage of the air conditioning and rising temperatures, we are forced to shut down the…

29.6.2019 - Restricted access to AV499 ("Tag der Offenen Tür")

On Saturday, June 29, the NMR lab in r.115 will be temporarily closed for user access

26.06.2019 AV499 out of Service

Dear User,

 

next Wednesday (26th of June) the av499 will be out of service for maintenance work.

 

It is expected to…

4.6.2019 - Nitrogen refills shifted

Next refills:

Tue, 11.06.2019

Tue, 18.06.2019

 

29.4.2019 - NMR-Abteilung: Important note for today!

For all of our spectrometers, a electric power bypass will be installed today. This work has not been announced to us…

16.4.2019 - NMR-Abteilung: OC-WP Praktikum schedule AV300oc

From 18th of April until 17th of May the walk-up mode is opened from 8:30am until 2pm. Afterwards it is switched into…

1.4.2019 - NMR-Abteilung: Limited access on April 10&11

On April 10th and 11th, several spectrometers will not be accessible for maintenance

26.3.2019 - NMR-Abteilung: Server Laplace down

The LIMS and spectra server is currently not operating

25.2.2019 - NMR-Abteilung: Karneval

The NMR facility remains cosed during carnival

zum Seitenanfang